OpenStudy (anonymous):

Given f(x) = -3x^2 + 4x, Find the following. f(a) f(x+1)

OpenStudy (yolo_king):

Degree = 2

OpenStudy (anonymous):

what do you mean by that? are you asking if that (^2) is "to the second degree" ?

OpenStudy (anonymous):

@YOLO_KING

OpenStudy (yolo_king):

Nvm I looked at the problem wrong

OpenStudy (yolo_king):

So your looking to factor it pretty much?

OpenStudy (yolo_king):

because \[f(x)=-3x^2+4x\]

OpenStudy (trojanpoem):

Not sure , f(a) = -3a^2 + 4 a f(x+1) = -3(x+1)^2 + 4(x+1) = -3(x^2 + 1 + 2x) + 4x + 4 = -3x^2 - 3 - 6x +4x + 4 -3x^2 - 2 x + 1 = f(x+1)

OpenStudy (yolo_king):

OOps i meant \[f(x) = -x(3x-4)\]

OpenStudy (anonymous):

Well the given part is f(x)=−3x2+4x on f(a) i got 3a^2 + 4(a) but i think theres another step

OpenStudy (anonymous):

@YOLO_KING @TrojanPoem

OpenStudy (trojanpoem):

you can do : f(a) = a(3a + 4)

OpenStudy (triciaal):

Given f(x) = -3x^2 + 4x, Find the following. f(a) f(x+1) to find f(a) replace x with a and rewrite to get f(a) = -3a^2 + 4a = a(4 - 3a) to find f(x + 1) replace x in f(x) with (x + 1 ) then simplify f(x + 1) = -3(x + 1)^2 + 4(x + 1) = (x +1)( 4 - 3(x + 1)) = (x + 1)(4 - 3x -3)= (x + 1)(1 - 3x) if you look closely you will see that after writing f(a) = a(4 - 3a) to find f(x + 1) you can let a = x + 1 and do the substitution